OpenStudy (anonymous):

Solve for x, given the equation Square root of x plus 9 − 4 = 1

OpenStudy (dehelloo):

sketchy as phuck The CEO of this website legit uploads pics of girls to gain internet traffic for lonesome fools to beat off to.

OpenStudy (jackthegreatest):

i thought it was teh square root of x

OpenStudy (anonymous):

Thanks, but here are the answer choices: x = 16, solution is not extraneous x = 16, solution is extraneous x = 34, solution is not extraneous x = 34, solution is extraneous

OpenStudy (anonymous):

 Is that it?

OpenStudy (anonymous):

You would subtract 9-4. 5. Square root of x + 5 = 1 Subtract 5. square root of x = -4 Square root is positive 2

OpenStudy (anonymous):

@Rishi_K Can you explain what you did? It make no sense....

OpenStudy (anonymous):

Step by step [Root of (x+9)]-4=1 [Root of (x+9)]=1+4 x+9=5^2 x=25-9 Therefore x=16

OpenStudy (anonymous):

@DIDITHEDAD